# A lower bound on the number of edges in DP-critical graphs

### v1

@@ 1.4 Outline of the proof and structure of the paper
    $\sum_{v\in V(G)}\left(w(v)-\sum_{e\in E(G):v\in e}\frac{w(e)}{2}\right)\leq-2.$ (8)
    Note that vertices of degree at most $k-1$ contribute positive terms to the sum, while vertices of degree at least $k$ contribute negative terms. Therefore, in order to prove the inequality, we plan to prove that the $k^{+}$-vertices of $G$ contribute enough to the sum in order to negate the positive terms from the vertices of degree at most $k-1$.
  3. 3.
    Sometimes it is convenient to consider smaller graphs obtained from $G$ in which some vertices have lists with fewer than $k-1$ colors. For induction, it is also convenient to consider loopless multigraphs rather than just simple graphs. Therefore, we develop a more general and sophisticated model in which the weights of vertices with smaller list sizes are less than $\lambda(k-1)+1$ and the weights of multiple edges are greater than $2\lambda$. In this new model, we still aim to prove (8).
  4. 4.
    In our new model, every positive term of (8) corresponds to a low vertex, that is, a vertex $v$ for which the list size $h(v)$ is equal to the degree $d_{G}(v)$. We let $\mathcal{B}$ be the set of low vertices in $G$, a minimum counterexample to our new model. We show that the induced subgraph $G[\mathcal{B}]$ is a GDP-forest.
  5. 5.
@@
    (a) We show that GDP-forests of maximum degree $k-1$ and without $(k-1)$-regular and $(k-2)$-regular blocks do not have too many edges.
    (b) We show that in a minimum counterexample $G$, $G[\mathcal{B}]$ has neither $(k-1)$-regular nor $(k-2)$-regular blocks.
@@
The structure of the paper is as follows. In Section 2, we introduce our model discussed above, which allows multiple edges and variable list sizes of the vertices. In Section 3 we discuss properties of GDP-forests. Properties of a minimum counterexample are shown in Section 4, and we finish the proof via the discharging method in Section 5.
2 The setup
@@
Let $G$ be a loopless multigraph, and let $h:V(G)\to\{0,\ldots,k-1\}$ be a function. For each $v\in V(G)$ and $xy\in{V(G)\choose 2}$, we define the potential $\rho_{G,h}(v)$ and $\rho_{G,h}(xy)$ to be as follows:
$\hfill\rho_{G,h}(v)=\begin{cases}h(v)\lambda+1&\textrm{if }h(v)=k-1\\ h(v)\lambda-1&\textrm{if }h(v)\in\{2,\ldots,k-2\}\\ h(v)\lambda-2&\textrm{if }h(v)\in\{0,1\}.\end{cases}\phantom{0}\hfill\phantom{0}\rho_{G,h}(xy)=\begin{cases}0&\textrm{if }xy\notin E(G)\\ 1-(2\lambda+1)|E_{G}(x,y)|&\textrm{otherwise.}\end{cases}\hfill$
In other words, if $x$ and $y$ are joined by a single edge in $G$, then $\rho_{G,h}(xy)=-2\lambda$, and each additional edge joining $x$ and $y$ adds $-(2\lambda+1)$ to $\rho_{G,h}(xy)$.
@@
Given a graph $G$, let $h:V(G)\to\mathbb{N}\cup\{0\}$. A cover $(H,L)$ of $G$ is an $h$-cover of $G$ if $|L(v)|\geq h(v)$ for each $v\in V(G)$. When $h\equiv t$ for some constant $t\in\mathbb{N}$, we call an $h$-cover simply a $t$-cover. We say that $G$ is DP $h$-colorable if $G$ has an $(H,L)$-coloring for every $h$-cover $(H,L)$ of $G$. We say that $G$ is (DP) $h$-minimal if $G$ is not DP $h$-colorable, but every proper subgraph $G^{\prime}$ of $G$ is DP $h|_{V(G^{\prime})}$-colorable. Given a vertex subset $A\subseteq V(G)$, we often write $h$ for the restriction $h|_{A}$ when this does not cause confusion.
@@
Let $k\geq 5$. Let $G$ be a loopless multigraph and $h:V(G)\to\{0,\ldots,k-1\}$. If $G$ is $h$-minimal, then one of the following holds:
  1. 1.
    $G=K_{k}$ and $h(v)=k-1$ for each $v\in V(G)$,
  2. 2.
    $k=5$, either $G=K^{4}_{2}$ or $G$ is a double cycle, and $h(v)=4$ for each $v\in V(G)$, or
  3. 3.
@@
Note that when $G$ is a simple graph and $h(v)=k-1$ for each $v\in V(G)$, the statement $\rho_{h}(G)\leq-2$ is equivalent to the statement $|E(G)|\geq(k-1+\frac{1}{\lambda})\frac{|V(G)|}{2}+\frac{1}{\lambda}$. Therefore, since $K^{4}_{2}$ and double cycles are not simple graphs, Theorem 2.1 implies Theorem 1.8.
A multigraph is exceptional if it is either a complete graph, or $K_{2}^{4}$ or a double cycle in the case that $k=5$.
We end this section with some lemmas for the potential function. The following well-known lemma (c.f. [19, Fact 11]) shows that our potential function is submodular.
@@
Let $F$ be a copy of $K_{k}$ and let $h(v)=k-1$ for each $v\in V(F)$. For a proper induced subgraph $F^{\prime}$ of $F^{-}$, $\rho_{h}(F^{\prime})\geq\rho_{h}(F^{-})+(k-3)(\lambda-1)-2$.
@@
The potential of a complete graph on $j$ vertices is
$j((k-1)\lambda+1)-2\lambda\binom{j}{2}=j\lambda k-j\lambda+j-\lambda j(j-1)=-\lambda j^{2}+j(\lambda k+1).$
In order to minimize the potential of $F^{\prime}$, we assume that $F^{\prime}$ is a complete graph. If $|V(F^{\prime})|=j$, then $\rho_{h}(F^{\prime})=-\lambda j^{2}+j(\lambda k+1)$. As a function of $j$, $\rho_{h}(F^{\prime})$ is a concave quadratic function and hence is minimized at $j=1$ or $j=k-1$. Therefore,
$\rho_{h}(F^{\prime})\geq\min\{(k-1)\lambda+1,(k-1)(\lambda+1)\}=(k-1)\lambda+1.$
Therefore, $\rho_{h}(F^{\prime})-\rho_{h}(F^{-})\geq(k-1)\lambda+1-(k+2\lambda)=(k-3)(\lambda-1)-2$. ∎
@@
As $k=5$, $\lambda=6$, so, $(k-1)\lambda+1=25$ and $122\lambda=$ 12. If $F=K_{5}$, then $\rho_{h}(F)=5(25)-10(12)=5$. If $F=K_{2}^{4}$, then $\rho_{h}(F)=2(25)-12-3(13)=-1$. If $F=C_{n}^{2}$, then $\rho_{h}(F)=25n-12n-13n=0$.
Let $A$ be a nonempty proper subset of $V(F^{-})$. We consider the following cases. If $F=K^{4}_{2}$, then $\rho_{h}(F^{-})=12$, and $\rho_{h}(A)=25$. If $F=K_{5}$, then $\rho_{h}(F^{-})=17$, and $\rho_{F^{-}}(A)\geq 25|A|-12\binom{|A|}{2}\geq 25$. If $F$ is a double cycle, then $\rho_{h}(F^{-})=13$ and $\rho_{F^{-},h}(A)\geq 25|A|-(12+13)(|A|-1)=25$.∎
3 Properties of GDP-forests
For a positive integer $q$, the $q$-blowup of a graph $G$ is the graph obtained by replacing each $v\in V(G)$ with an independent set $I_{v}$ of size $q$ and replacing each edge $uv\in E(G)$ with a $K_{q,q}$ joining $I_{u}$ and $I_{v}$.
Suppose that $(H,L)$ is a DP-cover of a multigraph $G$. The $q$-blowup of $(H,L)$ is a DP-cover of $G^{q}$ obtained by replacing each vertex $u\in V(H)$ with an independent set $I_{u}$ of size $q$, replacing each edge $uw\in E(H)$ with a $K_{q,q}$ joining $I_{u}$ and $I_{w}$, and replacing each list $L(v)$ with the set $\bigcup_{u\in L(v)}I_{u}$.
@@ 3 Properties of GDP-forests
Let $G$ be a loopless multigraph and $h$ be a function on $V(G)$. Let $(H,L)$ be an $h$-cover of $G$.
(a) If $G=C^{q}_{2t}$ and $h(v)=2q$ for all vertices $v$, then $G$ has no $(H,L)$-coloring if and only if $(H,L)$ is a $q$-blowup of a $2$-cover $(H^{\prime},L^{\prime})$ of $C_{2t}$ where $H^{\prime}$ is isomorphic to $C_{4t}$.
@@
(b) If $G=K^{q}_{t}$ and $h(v)=(t-1)q$ for all vertices $v$, then $G$ has no $(H,L)$-coloring if and only if $(H,L)$ is a $q$-blowup of a $(t-1)$-cover $(H^{\prime},L^{\prime})$ of $K_{t}$ where $H^{\prime}$ is isomorphic to $t-1$ disjoint copies of $K_{t}$.
@@
(c) If $G=C^{q}_{2t+1}$ and $h(v)=2q$ for all vertices $v$, then $G$ has no $(H,L)$-coloring if and only if $(H,L)$ is a $q$-blowup of a $2$-cover $(H^{\prime},L^{\prime})$ of $C_{2t+1}$ where $H^{\prime}$ is isomorphic to two disjoint copies of $C_{2t+1}$.
Proof.
In each case, the necessity direction is easy to check, so we prove sufficiency.
We first claim that in each case, $H$ is $d$-regular, where $d$ is the regularity of $G$. Indeed, suppose that $H$ contains a vertex $v$ with a color $c\in L(v)$ such that $d_{H}(c)<d$. Then, we assign $c$ to $v$. We write $G^{\prime}=G-v$, and for each $w\in V(G^{\prime})$, we write $L^{\prime}(w)$ for the set of colors in $L(w)$ that are not adjacent to $c$. Then, for each $w\in V(G^{\prime})$, $|L^{\prime}(w)|\geq d_{G^{\prime}}(w)$, and $|L^{\prime}(w^{*})|>d_{G^{\prime}}(w^{*})$ for at least one vertex $w^{*}\in V(G^{\prime})$. Therefore, by degeneracy, $G$ has an $(H,L)$-coloring that assigns $c$ to $v$, a contradiction. Therefore, we assume that $H$ is $d$-regular. In particular, for each adjacent pair $u,v\in V(G)$, $H[L(u)\cup L(v)]$ is a $q$-regular bipartite graph.
(a) First, we claim that $(H,L)$ is a $q$-blowup of some $2$-cover of $C_{2t}$. Consider two adjacent vertices $u,v\in V(G)$. By degeneracy, $G-\{u,v\}$ has an $(H,L)$-coloring $f$. For each $w\in\{u,v\}$, we write $L^{\prime}(w)$ for the subset of $L(w)$ consisting of colors with no neighbor in $f$. Note that $|L^{\prime}(w)|\geq q$ for each $w\in\{u,v\}$. If $H[L^{\prime}(u)\cup L^{\prime}(v)]$ is not a $K_{q,q}$, then $f$ extends to $u$ and $v$, a contradiction. Therefore, as $H[L(u)\cup L(v)]$ is $q$-regular, $H[L(u)\cup L(v)]$ consists of two disjoint copies of $K_{q,q}$.
Next, consider three consecutive vertices $u,v,w$ in $G$. By our previous observation, $H[L(u)\cup L(v)]$ consists of two disjoint copies $K_{uv}$ and $K_{uv}^{\prime}$ of $K_{q,q}$. Similarly, $H[L(v)\cup L(w)]$ consists of two disjoint copies $K_{vw}$ and $K_{vw}^{\prime}$ of $K_{q,q}$. Without loss of generality, $K_{uv}$ and $K_{vw}$ intersect in at least one vertex of $L(v)$. Now, suppose that $|V(K_{uv})\cap V(K_{vw})|<q$. Then, for each pair $(K^{\prime},K^{\prime\prime})$ such that $K^{\prime}\in\{K_{uv},K_{uv}^{\prime}\}$ and $K^{\prime\prime}\in\{K_{vw},K^{\prime}_{vw}\}$, we have $|V(K^{\prime})\cap V(K^{\prime\prime})|\in\{1,\ldots,q-1\}$. Let $f$ be an $(H,L)$-coloring of $G-v$. By our observation above, $f(u)$ and $f(w)$ have a common neighbor in $L(v)$. Therefore, at least one color in $L(v)$ is adjacent to neither $f(u)$ nor $f(w)$, and hence $f$ can be extended to $v$, a contradiction. Therefore, without loss of generality, $|V(K_{uv})\cap V(K_{vw})|=q$, and $|V(K^{\prime}_{uv})\cap V(K^{\prime}_{vw})|=q$. As this property holds for every consecutive triple $u,v,w$, it follows that $(H,L)$ is a $q$-blowup of a $2$-cover of $C_{2t}$.
Finally, up to isomorphism, there are only two $2$-covers $(H^{\prime},L^{\prime})$ of $C_{2t}$, namely one in which $H^{\prime}$ is isomorphic to two disjoint copies of $C_{2t}$ and one in which $H^{\prime}$ is isomorphic to $C_{4t}$. It is easy to check that if $H^{\prime}$ is isomorphic to two disjoint copies of $C_{2t}$, then $G$ admits an $(H,L)$-coloring. Therefore, $H^{\prime}$ is isomorphic to $C_{4t}$.
@@
(b) Suppose that $G=K^{q}_{t}$, $h(v)=(t-1)q$ for all vertices $v$, and $G$ has no $(H,L)$-coloring. We aim to show that $(H,L)$ is a $q$-blowup of a $(t-1)$-cover $(H^{\prime},L^{\prime})$ of $K_{t}$ where $H^{\prime}$ is isomorphic to $t-1$ disjoint copies of $K_{t}$.
We induct on $t$. If $t=1$, then as $h(v)=0$ for the single vertex $v\in V(G)$, $H$ is the empty graph, which is 00 disjoint $q$-blowups of $K_{1}$. Now, suppose that $t\geq 2$. Consider a vertex $v\in V(G)$, and write $G^{\prime}=G-v$. For each color $c\in L(v)$ and vertex $w\in V(G^{\prime})$, letting $L^{\prime}(w)$ be the set of colors in $L(w)$ that are not adjacent to $c$, $G^{\prime}$ admits no $(H-L(v),L^{\prime})$-coloring. Therefore, by the induction hypothesis, $H[L^{\prime}(V(G^{\prime}))]-L(v)$ is isomorphic to $t-2$ disjoint $q$-blowups of $K_{t-1}$. Thus, by the regularity of $H$, $H-L(v)$ is isomorphic to $t-1$ disjoint $q$-blowups of $K_{t-1}$. We write $A_{1},\dots,A_{t-1}$ for these $t-1$ disjoint $q$-blowups of $K_{t-1}$ that cover $G^{\prime}$.
Now, by our observation above, each color $c\in L(v)$ is adjacent to a single component $A_{i}$, and furthermore, for each $w\in V(G^{\prime})$, $c$ has a neighbor in all vertices of $L(w)\cap V(A_{i})$. Therefore, by the regularity of $H$, $H$ is isomorphic to $t-1$ disjoint $q$-blowups of $K_{t-1}$, and the result follows.
(c) If $t=1$, then we are in Case (b). If $t\geq 2$, then by following the same argument as in (a), $(H,L)$ is a $q$-blowup of a $2$-cover $(H^{\prime},L^{\prime})$ of $C_{2t+1}$. It is easy to check that if $H^{\prime}$ is isomorphic to $C_{4t+2}$, then $G$ has an $(H,L)$-coloring. Therefore, $H^{\prime}$ is isomorphic to two disjoint copies of $C_{2t+1}$. ∎
@@
For a multigraph $F$, let $\widetilde{F}$ denote the skeleton of $F$, i.e. the simple graph from which $F$ is obtained by multiplying some edges. In other words, $\widetilde{F}$ is the simple graph on $V(F)$ in which $u$ and $v$ are adjacent if and only if $|E_{F}(u,v)|\geq 1$. For a vertex $v\in V(F)$, let $\widetilde{d}(v)$ denote $d_{\widetilde{F}}(v)$, the degree of $v$ in $\widetilde{F}$, which is equal to $|N_{F}(v)|$.
For a multigraph $F$, let $m(F):=|E(F)|-|E(\widetilde{F})|$. So, $2m(F)=\sum_{v\in V(F)}(d_{F}(v)-\widetilde{d}_{F}(v))$.
@@
(iii) $T$ has neither $(k-1)$-regular nor $(k-2)$-regular blocks.
Then
$\Phi_{k}(T):=\alpha\sigma_{h}(T)+m(T)+|V^{-}_{k-1}(T)|-|V_{k-1}(T)|>1+\alpha.$ (9)
@@
for $k\geq 5$, again implying (9). Suppose $T$ has two vertices, say $v_{1}$ and $v_{2}$, and $j\geq 1$ edges, so $d_{T}(v_{1})=d_{T}(v_{2})=j$, $|V(T)|=2$ and $m(T)=j-1$. By (i)–(iii), $j\leq k-3$. We can write $\Phi_{k}(T)$ in the form $\Phi_{k}(T)=\sum_{i=1}^{2}\Phi^{\prime}_{k}(v_{i})$, where
@@
If $T$ has more than two vertices, then $T$ is a multiple of a cycle or a complete graph. In particular, $T$ is $j$-regular for some $31\leq j\leq k-$ 3. Similarly to above, we can write $\Phi_{k}(T)=m(T)+\sum_{v\in V(T)}\Phi_{k}(v)$, where
@@
$\Phi_{k}(v)=\alpha(k-1-j)+0-1\geq 3\alpha-1$ (10)
In this case, $\Phi_{k}(T)\geq|V(T)|\geq 3\cdot\min\{1,3\alpha-1\}>1+\alpha$, and we are done. Thus, we suppose $j=k-3$. Then instead of (10), we get
$\Phi_{k}(v)=\alpha(k-1-j)+0-1=2\alpha-1=\frac{3}{2k-7}>0.$ (11)
So if $|V_{k-1}^{-}(T)|+m(T)\geq 2$, then $\Phi_{k}(T)>2>1+\alpha$, as claimed. Otherwise, $G$ is a simple $(k-3)$-regular graph and $|V_{k-1}^{-}(T)|\leq 1.$ It follows that $G=K_{k-2}$ and by (11),
@@
Now, let $b\geq 1$, and suppose that the lemma holds for GDP-trees with at most $b$ blocks, and that $T$ has $b+1\geq 2$ blocks. Let $R$ be a root block in $T$, and let $B$ be a peripheral block, that is, a block in $T$ whose distance from $R$ in the block cut tree of $T$ is maximum. Choose $B$ so that the cut vertex $u$ in $B$ has the minimum number of neighbors in $B$. Let $s=|V(B)|-1$.
Let $T^{\prime}=T-(V(B)-u)$. Then $T^{\prime}$ is a GDP-tree satisfying (i)–(iii) and has fewer blocks than $T$. So, by induction, $\Phi_{k}(T^{\prime})>1+\alpha$. Since $T$ is a GDP-tree, $B$ is $j$-regular for some $31\leq j\leq k-$ 3. In particular, $d_{T}(u)-d_{T^{\prime}}(u)=j$, and the multiplicity of every existing edge in $B$ is the same, say $q$.

### v2

@@ 1.4. Outline of the proof and structure of the paper
    (9) $\sum_{v\in V(G)}\left(w(v)-\sum_{e\in E(G):v\sim e}\frac{w(e)}{2}\right)\leq-2.$
    Note that vertices of degree at most k1k-1 contribute positive terms to the sum, while vertices of degree at least $k$ contribute negative terms. Therefore, in order to prove the inequality, we plan to prove that the $k^{+}$-vertices of $G$ contribute enough to the sum in order to negate the positive terms from the vertices of degree at most k1k-1.
  3. (3)
    Sometimes it is convenient to consider smaller graphs obtained from $G$ in which some vertices have lists with fewer than k1k-1 colors. For induction, it is also convenient to consider multigraphs rather than just graphs. Therefore, we develop a more general and sophisticated model in which the weights of vertices with smaller list sizes are less than $\lambda(k-1)+1$ and the weights of multiple edges are greater than $2\lambda$. In this new model, we still aim to prove (9).
  4. (4)
    In our new model, every positive term of (9) corresponds to a low vertex, that is, a vertex $v$ for which the list size $h(v)$ is equal to the degree $d_{G}(v)$. We let $\Lambda$ be the set of low vertices in $G$, a minimum counterexample to our new model. We show that the induced subgraph $G[\Lambda]$ is a GDP-forest.
  5. (5)
@@
    (a) We show that GDP-forests of maximum degree k1k-1 and without (k1)(k-1)-regular and (k2)(k-2)-regular blocks do not have too many edges.
    (b) We show that in a minimum counterexample $G$, $G[\Lambda]$ has neither (k1)(k-1)-regular nor (k2)(k-2)-regular blocks.
@@
The structure of the paper is as follows. In Section 2, we introduce our model discussed above, which allows multiple edges and variable list sizes of the vertices. In Section 3 we discuss properties of GDP-forests. Properties of a minimum counterexample are shown in Section 4, and we finish the proof via the discharging method in Section 5. In Section 6 we prove the bound for $f_{\ell}(n,k)$. In conclusion, we briefly discuss possible further research directions.
2. The setup
@@
Let $G$ be a multigraph, and let $h:V(G)\to[0,k-1]$ be a function. For each $v,x,y\in V(G)$, we define the potential $\rho_{G,h}(v)$ and $\rho_{G,h}(xy)$ to be as follows:
$\hfill\rho_{G,h}(v)=\begin{cases}h(v)\lambda+1&\textrm{if }h(v)=k-1\\ h(v)\lambda-1&\textrm{if }h(v)\in[2,k-2]\\ h(v)\lambda-2&\textrm{if }h(v)\in\{0,1\}.\end{cases}\phantom{0}\hfill\phantom{0}\rho_{G,h}(xy)=\begin{cases}0&\textrm{if }|E_{G}(x,y)|=0\\ 1-(2\lambda+1)|E_{G}(x,y)|&\textrm{otherwise.}\end{cases}\hfill$
In other words, if $x$ and $y$ are joined by a single edge in $G$, then $\rho_{G,h}(xy)=-2\lambda$, and each additional edge joining $x$ and $y$ adds $(2\lambda+1)$ to $\rho_{G,h}(xy)$.
@@
Let $\mathbb{N}$ denote the set of positive integers. Given a multigraph $G$, let $h:V(G)\to\mathbb{N}\cup\{0\}$. A cover $(H,L)$ of $G$ is an $h$-cover of $G$ if $|L(v)|\geq h(v)$ for each $v\in V(G)$. When $h\equiv t$ for some constant $t\in\mathbb{N}$, we call an $h$-cover simply a $t$-cover. We say that $G$ is DP $h$-colorable if $G$ has an $(H,L)$-coloring for every $h$-cover $(H,L)$ of $G$. We say that $G$ is (DP) $h$-minimal if $G$ is not DP $h$-colorable, but every proper subgraph $G^{\prime}$ of $G$ is DP $h|_{V(G^{\prime})}$-colorable. Given a vertex subset $A\subseteq V(G)$, we often write $h$ for the restriction $h|_{A}$ when this does not cause confusion.
@@
Let $k\geq 5$. Let $G$ be a multigraph and $h:V(G)\to[0,k-1]$. If $G$ is $h$-minimal, then one of the following holds:
  1. (1)
    $G=K_{k}$ and $h(v)\in\{k-1,k-2\}$ for each $v\in V(G)$, with at most one vertex $w\in V(G)$ satisfying h(w)=k2h(w)=k-2,
  2. (2)
    $k=5$, either $G=4K_{2}$ or $G$ is a double cycle, and $h(v)=4$ for each $v\in V(G)$, or
  3. (3)
@@
Note that when $G$ is a graph and h(v)=k1h(v)=k-1 for each $v\in V(G)$, the statement $\rho_{h}(G)\leq-2$ is equivalent to the statement $|E(G)|\geq(k-1+\frac{1}{\lambda})\frac{|V(G)|}{2}+\frac{1}{\lambda}$. Therefore, since $4K_{2}$ and double cycles are not graphs, Theorem 2.1 implies Theorem 1.6.
A multigraph is exceptional if it is either a complete graph, or $4K_{2}$ or a double cycle in the case $k=5$.
We end this section with some lemmas for the potential function. The following well-known lemma (c.f. [26, Fact 11]) shows that our potential function is submodular.
@@
Let $G^{-}$ denote a multigraph obtained from $G$ by removing any edge.
@@
Let $F$ be a copy of $K_{k}$ and let $h(v)\in\{k-2,k-1\}$ for each $v\in V(F)$, where at most one vertex $v$ has h(v)=k2h(v)=k-2. For a proper induced subgraph $F^{\prime}$ of $F^{-}$, $\rho_{h}(F^{\prime})\geq\rho_{h}(F^{-})+(k-3)(\lambda-1)-2$.
@@
The potential of a complete graph $K_{j}$ having $j_{1}\in\{j-1,j\}$ vertices $v$ with h(v)=k1h(v)=k-1 and $j_{2}=j-j_{1}$ vertices $u$ with h(u)=k2h(u)=k-2 is
$j_{1}((k-1)\lambda+1)+j_{2}((k-2)\lambda-1)-2\lambda\binom{j}{2}=jk\lambda-j\lambda-j_{2}\lambda+j_{1}-j_{2}-\lambda j(j-1)$
$\lambda j^{2}+jk\lambda+j_{1}-j_{2}(\lambda+1).$
In order to minimize the potential of $F^{\prime}$, we assume that $F^{\prime}$ is a complete graph. If $|V(F^{\prime})|=j$, then $\rho_{h}(F^{\prime})=-\lambda j^{2}+j(\lambda k+1)$ when $j_{2}=0$ and $\rho_{h}(F^{\prime})=-\lambda j^{2}+j(\lambda k+1)-\lambda-2$ when $j_{2}=1$. In both cases, as a function of $j$, $\rho_{h}(F^{\prime})$ is a concave quadratic function and hence is minimized at $j=1$ or j=k1j=k-1. Therefore, if $j_{2}=0$, then
$\rho_{h}(F^{\prime})\geq\min\{(k-1)\lambda+1,(k-1)(\lambda+1)\}=(k-1)\lambda+1,$
and if $j_{2}=1$, then
$\rho_{h}(F^{\prime})\geq\min\{(k-2)\lambda-1,(k-2)(\lambda+1)-1\}=(k-2)\lambda-1.$
Thus, in both cases,
$\rho_{h}(F^{\prime})\geq(k-1)\lambda+1-j_{2}(\lambda+2).$
As $\rho_{h}(F^{-})=k-j_{2}(\lambda+2)+2\lambda$, we get
$\rho_{h}(F^{\prime})-\rho_{h}(F^{-})\geq(k-1)\lambda+1-k-2\lambda=(k-3)(\lambda-1)-2.$
@@
As $k=5$, $\lambda=6$, so, $(k-1)\lambda+1=25$ and $2\lambda=12$. If $F=K_{5}$, then $\rho_{h}(F)=5(25)-10(12)=5$. If $F=4K_{2}$, then $\rho_{h}(F)=2(25)-12-3(13)=-1$. If $F=2C_{n}$, then $\rho_{h}(F)=25n-12n-13n=0$.
Let $A$ be a nonempty proper subset of $V(F^{-})$. We consider the following cases. If $F=4K_{2}$, then $\rho_{h}(F^{-})=12$, and $\rho_{h}(A)=25$. If $F=K_{5}$, then $\rho_{h}(F^{-})=17$, and $\rho_{F^{-}}(A)\geq 25|A|-12\binom{|A|}{2}\geq 25$. If $F$ is a double cycle, then $\rho_{h}(F^{-})=13$ and $\rho_{F^{-},h}(A)\geq 25|A|-(12+13)(|A|-1)=25$.∎
3. Properties of GDP-forests
For a positive integer $q$, the $q$-blowup of a graph $G$ is the multigraph obtained by replacing each $v\in V(G)$ with an independent set $I_{v}$ of size $q$ and replacing each edge $uv\in E(G)$ with a $K_{q,q}$ joining $I_{u}$ and $I_{v}$.
Suppose that $(H,L)$ is a DP-cover of a multigraph $G$. The $q$-blowup of $(H,L)$ is a DP-cover of $qG$ obtained by replacing each vertex $u\in V(H)$ with an independent set $I_{u}$ of size $q$, replacing each edge $uw\in E(H)$ with a $K_{q,q}$ joining $I_{u}$ and $I_{w}$, and replacing each list $L(v)$ with the set $\bigcup_{u\in L(v)}I_{u}$.
The following lemma is a partial case of a result by Kim and Ozeki [21, Theorem 5], see also [31, Theorem 1.17].
@@ 3. Properties of GDP-forests
Let $G$ be a multigraph and $h$ be a function on $V(G)$. Let $(H,L)$ be an $h$-cover of $G$.
@@
(a) If $G=qK_{t}$ and h(v)=(t1)qh(v)=(t-1)q for all vertices $v$, then $G$ has no $(H,L)$-coloring if and only if $(H,L)$ is a $q$-blowup of a (t1)(t-1)-cover $(H^{\prime},L^{\prime})$ of $K_{t}$ where $H^{\prime}$ forms t1t-1 disjoint copies of $K_{t}$.
@@
(b) If $G=qC_{2t}$ and $h(v)=2q$ for all vertices $v$, then $G$ has no $(H,L)$-coloring if and only if $(H,L)$ is a $q$-blowup of a $2$-cover $(H^{\prime},L^{\prime})$ of $C_{2t}$ where $H^{\prime}$ is $C_{4t}$.
@@
(c) If $G=qC_{2t+1}$ and $h(v)=2q$ for all vertices $v$, then $G$ has no $(H,L)$-coloring if and only if $(H,L)$ is a $q$-blowup of a $2$-cover $(H^{\prime},L^{\prime})$ of $C_{2t+1}$ where $H^{\prime}$ forms two disjoint copies of $C_{2t+1}$.
For the next lemma, we need some more definitions and notation. Let ${\rm dist}_{G}(u,v)$ denote the distance from $u$ to $v$ in a multigraph $G$. The diameter, ${\rm diam}(G)$, of a connected multigraph $G$ is $\max_{u,v\in V(G)}{\rm dist}_{G}(u,v)$. A vertex $x\in V(G)$ is peripheral if there is $y\in V(G)$ such that ${\rm dist}_{G}(u,v)={\rm diam}(G)$. The block tree of a connected multigraph $G$ is the bipartite graph $H$ whose parts are the set of cut vertices and the set of the blocks of $G$, with a cut vertex $v$ adjacent to block $B$ in $H$ if and only if $v\in B$. This definition implies that the block tree of a connected multigraph $G$ is always a tree whose leaves are some blocks of $G$. Call a block of a connected multigraph $G$ peripheral if the corresponding vertex in the block tree of $G$ is peripheral.
@@
For a multigraph $F$, let $\widetilde{F}$ denote the underlying graph of $F$, i.e. the graph from which $F$ is obtained by multiplying some edges. In other words, $\widetilde{F}$ is the graph on $V(F)$ in which $u$ and $v$ are adjacent if and only if $|E_{F}(u,v)|\geq 1$. For a vertex $v\in V(F)$, let $\widetilde{d}(v)$ denote $d_{\widetilde{F}}(v)$, the degree of $v$ in $\widetilde{F}$, which is equal to $|N_{F}(v)|$.
For a multigraph $F$, let m(F):=|E(F)||E(F~)|m(F):=|E(F)|-|E($\widetilde${F})|. So, $2m(F)=\sum_{v\in V(F)}(d_{F}(v)-\widetilde{d}_{F}(v))$.
@@
(iii) $T$ has neither (k1)(k-1)-regular nor (k2)(k-2)-regular blocks.
Define $\Phi_{k}(T)=\alpha\sigma_{h}(T)+m(T)+|V^{-}_{k-1}(T)|-|V_{k-1}(T)|$. Then
(10) $\Phi_{k}(T)>1+\alpha.$
@@
for $k\geq 5$, again implying (10). Suppose $T$ has two vertices, say $v_{1}$ and $v_{2}$, and $j\geq 1$ edges, so $d_{T}(v_{1})=d_{T}(v_{2})=j$, $|V(T)|=2$ and m(T)=j1m(T)=j-1. By (i)–(iii), $j\leq k-3$. We can write $\Phi_{k}(T)$ in the form $\Phi_{k}(T)=\sum_{i=1}^{2}\Phi^{\prime}_{k}(v_{i})$, where
@@
If $T$ has more than two vertices, then $T$ is a multiple of a cycle or a complete graph. In particular, $T$ is $j$-regular for some $1\leq j\leq k-3$. Similarly to above, we can write $\Phi_{k}(T)=m(T)+\sum_{v\in V(T)}\Phi_{k}(v)$, where
@@
(11) $\Phi_{k}(v)=\alpha(k-1-j)+0-1\geq 3\alpha-1$
In this case, $\Phi_{k}(T)\geq|V(T)|\geq 3\cdot\min\{1,3\alpha-1\}>1+\alpha$, and we are done. Thus, we suppose j=k3j=k-3. Then instead of (11), we get
(12) $\Phi_{k}(v)=\alpha(k-1-j)+0-1=2\alpha-1=\frac{3}{2k-7}>0.$
So if $|V_{k-1}^{-}(T)|+m(T)\geq 2$, then $\Phi_{k}(T)>2>1+\alpha$, as claimed. Otherwise, $G$ is a (k3)(k-3)-regular graph and $|V_{k-1}^{-}(T)|\leq 1.$ It follows that $G=K_{k-2}$ and by (12),
@@
Now, let $b\geq 1$, and suppose that the lemma holds for GDP-trees with at most $b$ blocks, and that $T$ has $b+1\geq 2$ blocks. Choose a peripheral block $B$ in $T$ so that the cut vertex $u$ in $B$ has the minimum number of neighbors in $B$. Let s=|V(B)|1s=|V(B)|-1.
Let $T^{\prime}=T-(V(B)-u)$. Then $T^{\prime}$ is a GDP-tree satisfying (i)–(iii) and has fewer blocks than $T$. So, by induction, $\Phi_{k}(T^{\prime})>1+\alpha$. Since $T$ is a GDP-tree, $B$ is $j$-regular for some $1\leq j\leq k-3$. In particular, $d_{T}(u)-d_{T^{\prime}}(u)=j$, and the multiplicity of every existing edge in $B$ is the same, say $q$.